\begin{thebibliography}{99}

\setlength{\baselineskip}{15 pt}

\bibitem{BFMP2006} J. Bierbrauer, G. Faina, S. Marcugini and F. Pambianco, \emph{On the structure of the $(n,r)$-arcs in $PG(2,q)$}, Proceedings of the Tenth International Workshop on Algebraic and Combinatorial Coding Theory, Zvenigorod, Russia, 3-9 Settembre 2006, 19-23.

\bibitem{BKW2005} M. Braun, A. Kohnert and A. Wassermann, \emph{Construction of linear codes with prescribed distance}, OC05 The fourth International Workshop on Optima Codes and related Topics, PAMPOROVO, Bulgaria (2005), 59-63.

\bibitem{Cook2011} G. R. Cook, \emph{Arcs in a Finite Projective Plane}, PhD Thesis, available on internet http://sro.sussex.ac.uk/

%\bibitem{3ArchiPG2_13} K. Coolsaet, H. Sticker, \emph{The complete $(k,3)$-arcs of $PG(2,q)$, $q\leq13$}, Journal of Combinatorial Designs {\bf 20} (2012), 89-111.

\bibitem{H1988} J. W. P. Hirschfeld, \emph{Projective Geometries over Finite Fields}, second edition, Oxford University Press, Oxford, (1998).

\bibitem{HS2000} J. W. P. Hirschfeld and L. Storme, \emph{The packing problem in statistics, coding theory, and finite projective spaces}: update 2001, in: Finite Geometries, Proceedings of the Fourth Isle of Thorns Conference, A. Blokhuis, J. W. P. Hirschfeld, D. Jungnickel and J. A. Thas, Eds., Developments in Mathematics 3, Kluwer Academic Publishers, Boston, (2000), 201-246.

\bibitem{3ArchiPG2_7} S. Marcugini, A. Milani and  F. Pambianco, \emph{Classification of the $(n,3)$-arcs in $PG(2,7)$}, Journal of Geometry {\bf 80} (2004), 179-184.

\bibitem{3ArchiMPG2_11} S. Marcugini, A. Milani and F. Pambianco, \emph{Maximal $(n, 3)$-arcs in $PG(2, 11)$}, Proceedings of Discrete Mathematics (1999), 421-426.

\bibitem{3ArchiMPG2_13} S. Marcugini, A. Milani and F. Pambianco, \emph{Maximal $(n, 3)$-arcs in $PG(2, 13)$}, Proceedings of Discrete Mathematics (2005), 139-145.

\bibitem{Thas1975} J. Thas, \emph{Some results concerning $((q+1)(n-1),n)$-arcs}, J. Combin. Theory Ser. A {\bf 19} (1975), 228-232.

\bibitem{librogruppi}  A. D. Thomas and G. V. Wood, \emph{Group Tables}, Orpington, U.K.: Shiva mathematics series 2, (1980).

\end{thebibliography}

\documentclass[12pt,oneside, italian]{article}
\usepackage{latexsym}
\usepackage{color}
\usepackage{amssymb,amsthm,amsmath}
\usepackage{graphicx}

\newtheorem{theorem}{Theorem}

\newtheorem{conjecture}{Conjecture}
\begin{document}
\title{The maximum and the minimum size\\ of complete $(n,3)$-arcs in $PG(2,16)$}
\date{}
\author{D. Bartoli, S. Marcugini and F. Pambianco \\
%EndAName
Dipartimento di Matematica e Informatica,\\
Universit\`{a} degli Studi di Perugia, \\
Via Vanvitelli 1, 06123 Perugia Italy \\
e-mail: \{daniele.bartoli, gino, fernanda\}@dmi.unipg.it}
\maketitle
\setlength{\baselineskip}{15 pt}
\begin{abstract}
\noindent In this work we solve the packing problem for complete $(n,3)$-arcs in $PG(2,16)$, determining that the maximum size is $28$  and the minimum size is $15$. We also performed a partial classification of the extremal size of complete $(n,3)$-arcs in $PG(2,16)$.
\end{abstract}
%{\em Key words: }Covering codes, Projective geometry, Saturating sets

\section{Introduction}
In the projective plane $PG(2,q)$ over the finite field $GF(q)$ an $(n,r)$-arc is a set of
$n$ points such that no $r + 1$ points are collinear and some $r$ points are collinear. An $(n,r)$-arc is called \emph{complete} if it is not contained in a $(n+1,r)$-arc of the same projective plane. An $(n,2)$-arc is called $n$-arc. For a more detailed introduction to $(n,r)$-arcs and in particular $(n,3)$-arcs see \cite{H1988}, \cite{HS2000}, \cite{3ArchiMPG2_11},  \cite{3ArchiMPG2_13}. The largest size of $(n,r)$-arcs of $PG(2,q)$ is indicated by $m_{r}(2,q)$. In particular $m_{3}(2,q)\leq 2q+1$ for $q\geq 4$ (see \cite{Thas1975}). In \cite{HS2000} bounds for $m_{r}(2,q)$ and the relationship between the theory of complete $(n,r)$-arcs, coding theory and mathematical statistics are given.\\
Arcs and and $(n,3)$-arcs in $PG(2,q)$ correspond to respectively MDS and NMDS codes of dimension $3$. These types of linear codes are the best in term of minimum distance, among the linear codes with the same length and dimension. In general $(n,k)$-arcs in $PG(2,q)$ correspond to linear codes with Singleton defect equal to $k-2$.\\

\section{Results}
In this work we establish the maximum and the minimum size of complete $(n,3)$-arcs in $PG(2,16)$. To do this we performed a computer based search using some ideas similar to those presented in \cite{3ArchiPG2_7}, \cite{3ArchiMPG2_11}, \cite{3ArchiMPG2_13} and \cite{Cook2011}.
\begin{theorem}
The maximum size of complete $(n,3)$-arcs in $PG(2,16)$ is $28$.
\end{theorem}
\proof
We performed an exhaustive search of $(n,3)$-arcs in $PG(2,16)$ of size greater than $28$ and we found no examples. Moreover we have an example of complete $(28,3)$-arc (see \cite{BFMP2006}), obtained as union of orbits of some subgroup of $P\Gamma L(3,16)$; the size is equal to the one given in \cite{BKW2005}.
\endproof
The classification of complete $(28,3)$-arcs is in progress.

\begin{theorem}
The minimum size of complete $(n,3)$-arcs in $PG(2,16)$ is $15$.
\end{theorem}
\proof We performed an exhaustive search of $(n,3)$-arcs in $PG(2,16)$ of size less than
$15$ and we found no examples. We also proved that a complete $(15,3)$-arc contains a
$(k,2)$-arc, with $8\leq k \leq 9$. As result of the search for complete $(15,3)$-arcs we
get only the example presented in Table \ref{15arco} containing a $(9,2)$-arc.
\endproof

\begin{conjecture}
There exists a unique complete $(15,3)$-arc in $PG(2,16)$.
\end{conjecture}

We denote{\bf \ }$GF(16)=\{0,1=\alpha ^{0},2=\alpha ^{1},\ldots ,15=\alpha ^{14}\}$ where $%
\alpha $ is a primitive element such that $\alpha^4 + \alpha^3 + 1=0$. The columns $\ell_{i}$ indicate the number of $i$-secant of the $(n,3)$-arc and $G$ indicates the description of the stabilizer in $P\Gamma L(3,16)$ (see \cite{librogruppi}).

\begin{center}
\begin{table}
\caption{complete $(15,3)$-arc }\label{15arco}
\begin{center}
\begin{tabular}{|c|c|c|c|c|c|}
\hline
Points &$\ell_{0}$&$\ell_{1}$&$\ell_{2}$&$\ell_{3}$&G\\
\hline
\hline
\tabcolsep=0.85 mm
\begin{tabular}{ccccccccccccccc}
 1& 0& 0& 1& 1& 1& 1& 1& 1& 1& 1& 1& 1& 1& 1\\
 0& 1& 0& 1& 0& 1& 2& 2& 4& 9& 9& 11& 11& 13& 13\\
 0& 0& 1& 1& 11& 8& 5& 10& 10& 2& 8& 2& 11& 1& 12\\
\end{tabular}&
92 &
138 &
12 &
31 &
$\mathcal{S}_3$\\
\hline
\end{tabular}
\end{center}
\end{table}
\end{center}

\end{document}